function f(x) = 5/9(x-32) converts Fahrenheit temperatures into Celsius. What is the function for opposite conversion?
1
Expert's answer
2020-05-26T20:10:15-0400

Let "f(x)=y"

"\therefore" "y=\frac{5}{9}(x-32)"

"\implies 9y=5x-160 \implies9y+160=5x"

"\implies x=\frac{1}{5}(9y+160)" "=\frac{9}{5}y+32"

Therefore the opposite function ,which convert Celsius to Fahrenheit is "\frac{9}{5}y+32" ,where y is the temperature in Celsius. i,e,

"x°C,(\frac{9}{5}x+32)°F"